OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

find the slope first

OpenStudy (anonymous):

how do i find the slope ?

OpenStudy (anonymous):

Slope=y2−y1/x2−x1 Slope=2-1/-3-2=-1/5 So you have a slope of -1/5. You also have a given point. 2) Put into Point-Slope form. y−y1=m(x−x1) Lets use point (2, 1) and the slope of -1/5. try to solve:)

OpenStudy (anonymous):

follow @Aperogalics, I was trying to walk you through it

OpenStudy (anonymous):

so it would be D

OpenStudy (anonymous):

yup:)

OpenStudy (anonymous):

u got it congrats:)